# Action Reasoning Models that can Reason in Space

### v1

@@ 4.2 Mid-training
Following the initial pre-training stage, we conduct a second stage of mid-training using high-quality action reasoning data closely aligned with our target domain of household manipulation. Specifically, we formulate the MolmoAct Dataset into 1M action reasoning data samples and an additional 1M auxiliary trajectory-conditioned action data samples, which we find beneficial for improving overall performance and action steering. To train on the MolmoAct Dataset, we convert each sample—consisting of two side-mounted camera views and one wrist camera view, all sharing the same instruction and ground-truth action—into two paired-view training examples by pairing each side view with the wrist view. For action reasoning data preparation (subsection 3.1), depth perception tokens and visual reasoning traces are generated only from the side views, while the wrist view is used solely for action grounding. We fine-tune the model on this augmented dataset for 50k gradient steps with a batch size of 128, using 128 H100 GPUs over 2,304 GPU hours. Additional hyperparameters and optimization details are provided in Appendix B.
@@ 4.3 Post-training
After mid-training, we conduct the final post-training stage to rapidly adapt the model to new tasks and embodiments. For new tasks, we collect a small set of 30 to 50 tele-operated demonstrations per task, then generate perception tokens and visual reasoning trace tokens for each timestep. These demonstrations are converted into the action reasoning data and trajectory-conditioned action data, following the same process as mid-training data curation, with one key difference: to enhance manipulation performance and prevent no-ops from degrading results, we apply Action-Chunking (Zhao et al., 2023) during post-training, formatting action predictions in fixed-length chunks ($N=8$). We format the action chunks in the same way as we do for single action. After grouping them to a list, we train the model autoregressively for predicting all action chunks. We adapt MolmoAct to fine-tune downstream single or multi-tasks via parameter-efficient LoRA fine-tuning. In all evaluations, we fix the LoRA rank at 32 and alpha at 16 to preserve the model’s pre-trained capabilities. For simulation benchmarks (e.g., LIBERO), we use a batch size of 128, and for real-world tasks we use 64. The number of gradient steps we train varies by task. Post-training data generally consists of a front- or side-view image paired with a wrist-view image, although some setups provide multiple wrist views (e.g., bimanual scenarios). In every case, we apply the same LoRA and training configuration described above. Additional details are available in Appendix B.
@@ 5 Experimental Evaluation
Our experimental evaluation comprises a broad suite of studies that rigorously benchmark MolmoAct against strong baselines. We assess MolmoAct with MolmoAct-7B-D version in (i) its pre-training, “out-of-the-box” capabilities, (ii) its post-training adaptability across varied tasks, domains, and robotic embodiments, and MolmoAct-7B-D-Pretrain for (iii) its additional feature of being an interactive and steerable action reasoning model. By testing the model on a comprehensive range of scenarios both in simulation and real-world, we aim to answer the following research questions:
@@
    How effectively does MolmoAct follow language commands? We benchmark MolmoAct against strong baselines in an open-ended simulation setup where human evaluators issue free-form prompts and assess each model’s resulting actions.
@@
5.1 Evaluation of MolmoAct pre-trained model
@@
Evaluation Setup and Baselines. We first evaluated MolmoAct zero-shot—immediately after pre-training and before any task-specific fine-tuning. Unlike $\pi_{0}$ (Black et al., ), GR00T N1 (NVIDIA et al., 2025), and other proprietary VLA models that rely on large-scale private robot datasets and the full OXE dataset for pre-training, MolmoAct was trained exclusively on curated action reasoning data filtered from a subset of OXE (specifically BC-Z (Jang et al., 2022), BridgeData V2 (Walke et al., 2023), and RT-1 (Brohan et al., 2022)), combined with embodied reasoning VQA and auxiliary robot data, as detailed in Section 4.1. This amounts to approximately 26.3M data points—an order of magnitude smaller than $\pi_{0}$, which uses at least 903M for pre-training. To evaluate MolmoAct’s out-of-the-box generalization, we used the SimplerEnv benchmark, which features both visual-matching and variant-aggregation tasks across WidowX and Google Robot platforms. As MolmoAct ’s pre-training distribution is most aligned with the Google Robot visual-matching tasks, we focused our evaluation on this suite to best isolate in-distribution performance and capabilities of pre-training.
@@
We compared MolmoAct against a set of generalist policies, including TraceVLA (Zheng et al., 2024), RT-1X (Brohan et al., 2022), OpenVLA (Kim et al., 2024), RoboVLM (Liu et al., 2025), Emma-x (Sun et al., 2024), $\pi_{0}$ and $\pi_{0}$-FAST (Black et al., ), Octo (Team et al., 2024b), Magma (Yang et al., 2025b), HPT (Wang et al., 2024a), SpatialVLA (Qu et al., 2025) and GR00T N1 (NVIDIA et al., 2025). Most baselines were evaluated in the zero-shot setting, with a subset also tested after fine-tuning. We additionally fine-tuned MolmoAct on the RT-1 subset of OXE to assess its capacity when given more pre-training data.
@@
Evaluation Results. MolmoAct achieved strong zero-shot performance on the SimplerEnv visual-matching suite, reaching 70.5% accuracy and outperforming baselines such as GR00T N1, $\pi_{0}$, $\pi_{0}$-FAST, and Magma. With fine-tuning on the same RT-1 subset of OXE, MolmoAct-7B-D improved to 71.6%, exceeding Magma by 3.2% as shown in Table 1. These results indicate that MolmoAct is both an effective zero-shot generalist and a strong initialization for fine-tuned deployment.
5.2 Efficiency of MolmoAct fine-tuning
Evaluation Setups and Baselines. We evaluate MolmoAct in both simulation and real-world settings to assess its generalization capabilities after rapid fine-tuning. In simulation, we follow the OpenVLA evaluation protocol and benchmark MolmoAct on the LIBERO simulation benchmark, which includes diverse robotic manipulation tasks with 50 human-teleoperated demonstrations per task. We consider four subsets targeting different generalization axes: LIBERO-Spatial (spatial reasoning), LIBERO-Object (object variation), LIBERO-Goal (task abstraction), and LIBERO-Long (long-horizon planning). MolmoAct is fine-tuned using Low-Rank Adaptation (LoRA) and compared to state-of-the-art generalist policies that are autoregressive in nature, including TraceVLA (Zheng et al., 2024), OpenVLA (Kim et al., 2024), SpatialVLA (Qu et al., 2025), $\pi_{0}$-FAST (Black et al., ), CoT-VLA (Zhao et al., 2025), WorldVLA (Cen et al., 2025), ThinkAct (Huang et al., 2025), and NORA-AC (Hung et al., 2025). In the real world, we evaluate MolmoAct on six tasks across single-arm and bimanual Franka setups. The single-arm tasks include Put_Bowl_in_Sink, Wipe_Table, and Table_Bussing. The bimanual tasks include Set_Table, Lift_Box, and Fold_Towel. Each task is paired with 50 human demonstrations. We fine-tune both MolmoAct and baseline models on this data and evaluate performance over 25 trials per task using the task progression metric described in Appendix D. For both simulation and real-world evaluations, we used the MolmoAct model after it was mid-trained on our MolmoAct Dataset for LoRA fine-tuning. This setup enables a comprehensive comparison of adaptation efficiency across tasks and embodiments.
Evaluation Results.
@@ 5.2 Efficiency of MolmoAct fine-tuning
We evaluate on the LIBERO simulation benchmark (Liu et al. (2023a)), which consists of a Franka Emika Panda arm in simulation with demonstrations containing front and wrist view camera images ($\times256$ px), tasks language instructions, and delta end-effector pose actions. We follow prior works (Kim et al. (2024)) and evaluate on the four task suites – LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long – each with 500 demonstration across 10 tasks. Following (Kim et al. (2024)), we trained on a modified dataset which filtered out no-ops actions and unsuccessful demonstrations. Moreover, we set action chunk size to $K=8$ for evaluation on each task suites and execute full chunks before redoing action reasoning.
@@
On the LIBERO benchmark, MolmoAct achieves an average success rate of 86.6%, the highest among all compared methods. It performs particularly well on LIBERO-Long, a challenging long-horizon suite, where it exceeds the performance of ThinkAct—the second-best method in this setting—by 6.3%. MolmoAct also scores highest on LIBERO-Goal, achieving 87.6% as shown in Table 2, indicating strong generalization to task-oriented behaviors. In the real world, MolmoAct demonstrates effective fine-tuning and generalization across different embodiments. It outperforms $\pi_{0}$-FAST by an average of 10% in task progression on single-arm tasks and by 22.7% on bimanual tasks depicted in Figure 5.
@@
Figure 7: Line steerability evaluation across models. Left: Elo ratings show that MolmoAct achieves the highest performance, surpassing Gemini-2.5-Flash, GPT-4o, and HAMSTER, with error bars indicating standard error. Right: Example qualitative results showing predicted visual traces overlaid on robot camera views.
@@ 5.3 Effectiveness of MolmoAct in Out-of-Distribution Generalization
We evaluate MolmoAct in both simulation and real-world settings to assess its ability to generalize beyond the training data distribution, both in zero-shot and fine-tuned regimes. In simulation, we follow the SimplerEnv variant-aggregation protocol, which introduces distribution shifts through changes in lighting, textures, and camera viewpoints. We compare MolmoAct against several state-of-the-art generalist policies—TraceVLA, RT-1X, OpenVLA, RoboVLM, Emma-X, $\pi_{0}$-FAST, and SpatialVAL. For real-world evaluation, we test MolmoAct using a single Franka arm on a multi-task setup involving three objects and two different-colored plates arranged on a tabletop. We collect over 300 demonstrations spanning three task types, then train MolmoAct and baselines in a multi-task setting. During evaluation, we test generalization across four axes: (1) Language Variation — rephrased instructions, (2) Spatial Variation — changes in target object position, (3) Distractors — addition of irrelevant objects, and (4) Novel Objects — substitution of target objects with unseen ones. We benchmark MolmoAct against $\pi_{0}$-FAST and OpenVLA, testing three variants per task and four trials per variant.
Evaluation Results. In simulation, fine-tuned MolmoAct achieves 72.1% on the variant aggregation tasks as shown in Table 1, outperforming all baselines and exceeding the second-best model, RT-2-X, by 7.8%. The performance difference between variant aggregation and visual matching is less than 1%, highlighting MolmoAct’s robustness to visual and distributional shifts. In the real world, MolmoAct consistently surpasses all baselines across all generalization axes, achieving a 23.3% average improvement in task progression over $\pi_{0}$-FAST as shown in Figure 6(a).
@@ 5.4 Effect of the MolmoAct Dataset on MolmoAct performance
Evaluation Setups and Baselines. To assess the effectiveness of mid-training with the MolmoAct Dataset, we conducted real-world experiments on three curated tasks that goes beyond simple pick-and-place: close_lid, rotate_pot, and pour_tea. For each task, we collected 50 demonstrations and trained four models: MolmoAct-7B-D, MolmoAct-7B-D without mid-training, $\pi_{0}$-FAST, and OpenVLA. Each model was then evaluated over 10 trials per task.
Evaluation Results. Based on the real-world ablation studies shown in Figure 6(b), MolmoAct outperforms its counterpart without mid-training by an average margin of 5.5% across the three tasks, demonstrating that mid-training on the MolmoAct Dataset yields a consistent performance boost of around 5%. Even without mid-training, MolmoAct surpasses $\pi_{0}$-FAST and OpenVLA by 14.8% and 10.9%, respectively.
@@ 5.5 Instruction following of MolmoAct
We evaluated MolmoAct’s ability to follow natural language instructions in two settings: (i) executing open-ended commands in simulation, and (ii) generating visual traces conditioned on language prompts. For simulation-based instruction following, we curated five manipulation scenarios in the SimplerEnv environment using a Google Robot, each involving novel out-of-distribution objects. Ten participants provided 29 open-ended instructions (e.g., “Put the redbull into the bowl."). We compared MolmoAct to SpatialVLA and OpenVLA, both pretrained on the OXE dataset. For each instruction, the models generated trajectory rollouts, which were evaluated in a head-to-head arena-style web interface. Human annotators (n=100n{=}100) selected which trajectory best matched the instruction. We collected over 1,500 votes, which were converted into Elo ratings (see Figure 8). For visual trace generation, ten participants wrote 87 language prompts grounded in 30 internet-sourced images depicting tabletop and mobile manipulation contexts. MolmoAct was evaluated against Gemini-2.5-Flash, GPT-4o, and HAMSTER—a VLM fine-tuned for trace generation. Participants voted in a similar blind arena interface, resulting in over 1,000 votes.
Evaluation Results. MolmoAct achieved the highest Elo rating in the simulation instruction-following task, outperforming SpatialVLA by 109 points and OpenVLA by an even larger margin. Pairwise win rates also shown that MolmoAct winning over SpatialVLA in 58% of comparisons and over OpenVLA in 81%. A sample trajectory comparison for the instruction “Put the redbull into the bowl." is shown on the right. In the visual trace task, MolmoAct again outperformed all baselines, achieving significantly higher Elo scores with non-overlapping 95% confidence intervals, demonstrating strong generalization in both spatial execution and language-grounded reasoning.

### v2

@@ 4.2 Mid-training
Following the initial pre-training stage, we conduct a second stage of mid-training using high-quality action reasoning data closely aligned with our target domain of household manipulation. Specifically, we formulate the MolmoAct Dataset into 1M action reasoning data samples and an additional 1M trajectory-conditioned action data samples, which we find beneficial for improving overall performance and action steering. To train on the MolmoAct Dataset, we convert each sample—consisting of two side-mounted camera views and one wrist camera view, all sharing the same instruction and ground-truth action—into two paired-view training examples by pairing each side view with the wrist view. For action reasoning data preparation (subsection 3.1), depth perception tokens and visual reasoning traces are generated only from the side views, while the wrist view is used solely for providing additional information. We train the model on this modified dataset for 50k gradient steps with a batch size of 128, using 128 H100 GPUs over around 2,304 GPU hours. Additional hyperparameters and optimization details are provided in Appendix B.
@@ 4.3 Post-training
After mid-training, we conduct the final post-training stage to rapidly adapt the model to new tasks and embodiments. For new tasks, we collect a small set of 30 to 50 tele-operated demonstrations per task, then generate perception tokens and visual reasoning trace for each timestep. These demonstrations are converted into the action reasoning data and trajectory-conditioned action data, following the same process as mid-training data curation, with one key difference: we apply action chunking (Zhao et al., 2023) during post-training, formatting action predictions in fixed-length chunks ($N=8$). For each action chunk in all chunks, we tokenize each of them in the same way as we do for single action. After grouping them to a list, we train the model autoregressively for predicting all action chunks. We adapt MolmoAct to post-train on single or multi-tasks via parameter-efficient LoRA fine-tuning. In all evaluations, we fix the LoRA rank at 32 and alpha at 16 to preserve the model’s pre-trained capabilities. For simulation benchmarks (e.g., LIBERO), we use a batch size of 128, and for real-world tasks we use 64. The number of gradient steps we train varies by task. Our post-training data generally consists of a front- or side-view image paired with a wrist-view image, although some setups provide multiple wrist views (e.g., bimanual scenarios). In all cases, we apply the same LoRA and training configuration described above. Additional details are available in Appendix B.
@@ 5 Experimental Evaluation
Our experimental evaluation comprises a broad suite of studies that rigorously benchmark MolmoAct against strong baselines. We assess MolmoAct with MolmoAct-7B-D version in (i) its pre-training, “out-of-the-box” capabilities, (ii) its post-training adaptability across varied tasks, domains, and robotic embodiments, and (iii) its additional feature of being an interactive and steerable action reasoning model. By testing the model on a comprehensive range of scenarios both in simulation and real-world, we aim to answer the following research questions:
@@
    How effectively does MolmoAct follow language commands? We benchmark MolmoAct against strong baselines in an open-ended simulation setup where human evaluators provide free-form prompts and assess each model’s resulting actions.
@@
5.1 MolmoAct After Pre-training
@@
Evaluation Setup and Baselines. We first evaluated MolmoAct’s zero-shot capability—its ability immediately after pre-training and before any task-specific fine-tuning. Unlike $\pi_{0}$ (Black et al., ), GR00T N1 (NVIDIA et al., 2025), and other proprietary VLA models that rely on large-scale private robot datasets and the full OXE dataset for pre-training, MolmoAct was trained exclusively on curated action reasoning data filtered from a subset of OXE (specifically BC-Z (Jang et al., 2022), BridgeData V2 (Walke et al., 2023), and RT-1 (Brohan et al., 2022)), combined with multimodal web data and auxiliary robot data, as detailed in subsection 4.1. This amounts to approximately 26.3M samples—an order of magnitude smaller than $\pi_{0}$, which uses at least 903M for pre-training. To evaluate MolmoAct’s out-of-the-box generalization, we used the SimplerEnv benchmark, which features both visual-matching and variant-aggregation tasks across WidowX and Google Robot platforms (details in subsection D.1). As MolmoAct ’s pre-training distribution is most aligned with the Google Robot visual-matching tasks, we focused our evaluation on this suite to best isolate in-distribution performance and capabilities of pre-training.
@@
We compared MolmoAct-7B-D-Pretrain against a set of generalist policies, including TraceVLA (Zheng et al., 2024), RT-1X (Brohan et al., 2022), OpenVLA (Kim et al., 2024), RoboVLM (Liu et al., 2025), Emma-x (Sun et al., 2024), $\pi_{0}$ and $\pi_{0}$-FAST (Black et al., ), Octo (Team et al., 2024b), Magma (Yang et al., 2025b), HPT (Wang et al., 2024a), SpatialVLA (Qu et al., 2025) and GR00T N1 (NVIDIA et al., 2025). Most baselines were evaluated in the zero-shot setting, with a subset also tested after fine-tuning. We additionally fine-tuned MolmoAct-7B-D-Pretrain on the RT-1 subset of OXE to assess its capacity when given more pre-training data.
@@
Evaluation Results. MolmoAct-7B-D-Pretrain achieved strong zero-shot performance on the SimplerEnv visual-matching suite, reaching 70.5% success rate and outperforming baselines such as GR00T N1, $\pi_{0}$, $\pi_{0}$-FAST, and Magma. With fine-tuning on the same RT-1 subset of OXE, MolmoAct-7B-D improved to 71.6%, exceeding Magma by 3.2% as shown in Table 1. These results indicate that MolmoAct is both an effective zero-shot generalist and a strong initialization for fine-tuned deployment.
5.2 Fast Adaptation of MolmoAct in Post-training
@@
Evaluation Setups and Baselines. We evaluate MolmoAct in both simulation and real-world settings to assess its fast adaptation after post-training. In simulation, we evaluate on the LIBERO simulation benchmark (Liu et al. (2023a)), which consists of a Franka Emika Panda arm in simulation with demonstrations containing front and wrist view camera images ($\times256$ px), language instructions, and delta end-effector pose actions. We follow prior works (Kim et al. (2024)) and evaluate on four task suites – LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long – each with 500 demonstrations across 10 tasks. Following (Kim et al. (2024)), we trained on a modified dataset that filtered out no-op actions and unsuccessful demonstrations. Moreover, we set action chunk size to $K=8$ for evaluation on each task suite and execute full chunks before redoing action reasoning. We fine-tune MolmoAct-7B-D using Low-Rank Adaptation (LoRA) and compared to state-of-the-art generalist autoregressive policies, such as TraceVLA (Zheng et al., 2024), OpenVLA (Kim et al., 2024), SpatialVLA (Qu et al., 2025), $\pi_{0}$-FAST (Black et al., ), CoT-VLA (Zhao et al., 2025), WorldVLA (Cen et al., 2025), ThinkAct (Huang et al., 2025), and NORA-AC (Hung et al., 2025).
In the real world, we evaluate MolmoAct on six tasks across single-arm and bimanual Franka setups. The single-arm tasks include put_bowl_in_sink, wipe_table, and table_bussing. The bimanual tasks include set_table, lift_tray, and fold_towel. For each task, we collected 50 human tele-operated demonstrations and post-trained both MolmoAct-7B-D and baseline models. We evaluate the task progress over 25 trials per task. We detail the task progress scores in subsection D.3. This setup enables a comprehensive comparison of adaptation efficiency across tasks and embodiments.
@@
Evaluation Results. On the LIBERO benchmark, MolmoAct-7B-D achieves an average success rate of 86.6%, the highest among all compared methods. It performs particularly well on LIBERO-Long, a challenging long-horizon suite, where it exceeds the performance of ThinkAct—the second-best method in this setting—by 6.3%. In the real world, MolmoAct demonstrates effective fine-tuning and generalization across different embodiments. It outperforms $\pi_{0}$-FAST by an average of 10% in task progression on single-arm tasks and by 22.7% on bimanual tasks, as shown in Figure 5.
@@
Figure 7: Line steerability evaluation across models. Left: Elo ratings show that MolmoAct achieves the highest performance, surpassing Gemini-2.5-Flash, GPT-4o, and HAMSTER, with error bars indicating 95% confidence interval (CI). Right: Example qualitative results showing predicted visual traces overlaid on robot camera views.
@@ 5.3 Effectiveness of MolmoAct in Out-of-Distribution Generalization
We evaluate MolmoAct in both simulation and real-world settings to assess its ability to generalize beyond the training data distribution, both in zero-shot and fine-tuned regimes. In simulation, we follow the SimplerEnv variant-aggregation protocol, which introduces distribution shifts through changes in lighting, textures, and camera viewpoints. We compare MolmoAct-7B-D-Pretrain and its RT-1 fine-tuned variant against several state-of-the-art generalist policies—TraceVLA, RT-1X, OpenVLA, RoboVLM, Emma-X, $\pi_{0}$-FAST, and SpatialVLA. For real-world evaluation, we test MolmoAct-7B-D using a single Franka arm on a multi-task setup involving three objects and two different-colored plates arranged on a tabletop. We collect over 300 tele-operated demonstrations spanning three task types, then post-train MolmoAct-7B-D and baselines in a multi-task setting. During evaluation, we test generalization in four aspects: (1) Language Variation — rephrased instructions, (2) Spatial Variation — changes in target object position, (3) Distractors — addition of irrelevant objects, and (4) Novel Objects — substitution of target objects with unseen ones. We benchmark MolmoAct-7B-D against $\pi_{0}$-FAST and OpenVLA, testing three variants per task and four trials per variant. Full task details are presented in subsection D.4.
Evaluation Results. In simulation, fine-tuned MolmoAct-7B-D-Pretrain achieves 72.1% on the variant aggregation tasks as shown in Table 1, outperforming all baselines and exceeding the second-best model, RT-2-X, by 7.8%. The performance difference between variant aggregation and visual matching is less than 1%, highlighting MolmoAct’s robustness to visual and distributional shifts. In the real world, MolmoAct-7B-D consistently surpasses all baselines across all generalization axes, achieving a 23.3% average improvement in task progression over $\pi_{0}$-FAST as shown in Figure 6(a).
@@ 5.4 Effect of the MolmoAct Dataset on MolmoAct Performance
Evaluation Setups and Baselines. To assess the effectiveness of mid-training with the MolmoAct Dataset, we conducted real-world experiments on three curated tasks that go beyond simple pick-and-place: close_lid, rotate_pot, and pour_tea. For each task, we collected 50 demonstrations and trained four models: MolmoAct-7B-D, MolmoAct-7B-D without mid-training, $\pi_{0}$-FAST, and OpenVLA. Each model was then evaluated over 10 trials per task. Task details are shown in subsection D.5.
Evaluation Results. Based on the real-world ablation studies shown in Figure 6(b), MolmoAct-7B-D outperforms its counterpart without mid-training by an average margin of 5.5% across the three tasks, demonstrating that mid-training on the MolmoAct Dataset yields a consistent performance boost of around 5%. Even without mid-training, MolmoAct-7B-D-Pretrain surpasses $\pi_{0}$-FAST and OpenVLA by 14.8% and 10.9%, respectively.
@@ 5.5 Instruction Following of MolmoAct
We evaluated MolmoAct’s ability to follow natural language instructions in two settings: (i) executing tasks with open-ended commands in simulation, and (ii) generating visual traces conditioned on language prompts. For the first one, we curated five manipulation scenarios in the SimplerEnv environment using a Google Robot, each involving novel out-of-distribution objects. Ten participants provided 29 open-ended instructions (e.g., “Put the redbull into the bowl."). We compared MolmoAct-7B-D-Pretrain to SpatialVLA and OpenVLA, both pre-trained on the OXE dataset. For each instruction, the models generated rollouts, which were evaluated in a head-to-head arena-style web interface. Human annotators (n=100n{=}100) selected which rollout best matched the instruction. We collected over 1,500 votes, which were converted into Elo ratings (see Figure 8). For visual trace generation, 10 participants wrote 87 language prompts for 30 internet-sourced images depicting tabletop and mobile manipulation scenarios. MolmoAct-7B-D-Pretrain was evaluated against Gemini-2.5-Flash, GPT-4o, and HAMSTER—a VLM fine-tuned for trace generation. Participants voted in a similar blind arena interface, resulting in over 1,000 votes. Details with our curated manipulation scenes and instructions provided by participants are shown in subsection D.6.
Evaluation Results. MolmoAct-7B-D-Pretrain achieved the highest Elo rating in the simulation instruction-following task, outperforming SpatialVLA by 109 points and OpenVLA by an even larger margin. Pairwise win rates also show that MolmoAct-7B-D-Pretrain winning over SpatialVLA in 58% of comparisons and over OpenVLA in 81%. A sample rollout comparison for the instruction “Put the redbull into the bowl." is shown on the right in Figure 8. In the visual trace task, MolmoAct-7B-D-Pretrain again outperformed all baselines, achieving significantly higher Elo scores with non-overlapping 95% confidence intervals, demonstrating strong language-grounded generalization in both action execution and trace generation as shown in Figure 7.